What mass of HCIO4 must be present in 0.500 L of solution to obtain a solution with a pH=3.5 value
1
Expert's answer
2020-05-08T14:10:32-0400

Solution.

"pH = -lg[H3O^+]"

"[H3O^+] = 10^{-pH}"

"[H3O^+] = 3.16 \times 10^{-4} M"

"C = \frac{m}{M \times V}"

"m = C \times M \times V"

"m = 3.16 \times 10^{-4} \times 100.46 \times 0.500 = 0.016 \ g"

Answer:

m(HClO4) = 0.016 g